### v1

@@ B.4 Comparing Adaptation Time & Annotation Time
We carried out empirical studies by measuring the wall-clock time lapses of the inference step (Inference Time), adaptation step (Adaptation Time), and human annotation (Annotation Time) per sample. In specific, the "Inference Time" refers to the time-lapse of pure forward pass, the "Adaption Time" refers to the time-lapse of backpropagation and weights update, and "Annotation Time" refers to the estimated time lapse required for human to annotate images at a 3% labeling rate. We evaluate under three categories of annotation cost, i.e. 2 sec/sample, 5 sec/sample, and 10 sec/sample. As seen from Tab. 8, despite the adaptation time being higher than the inference time (8.7 ms v.s. 0.4 ms), the annotation time is the most expensive step. These results suggest model selection and adaptation do not slow down the throughput under the HILTTA protocol.
@@
B.5 Computation Complexity of Competing Methods
We conclude that existing stream-based active learning (VeSSAL [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth]) and active TTA ( SimATTA [Gui et al.(2024)Gui, Li, and Ji]) are inherently more computationally expensive than HILTTA due to the following reasons. First, VeSSAL adopts gradient embedding, i.e. $g(x_{i})=\partial l(f(x_{i};\theta),y_{i})/\partial\theta_{L}$, where $\theta_{L}\in\mathbb{R}^{K\times D}$ refers to the classifier weights. This results in a very high dimensional gradient embedding $g(x_{i})\in\mathbb{R}^{D\cdot K}$. The probability of labeling a sample $p_{i}$ involves calculating the inverse of covariance matrix, i.e. $p_{i}\propto g(x_{i})^{\top}\hat{\Sigma}^{-1}g(x_{i}),\;s.t.\hat{\Sigma}=\sum_{x_{i}\in\mathcal{B}}g(x_{i})g(x_{i})^{\top}\in\mathbb{R}^{DK\times DK}$. Hence, VeSSAL can hardly deal with classification tasks with a large number of classes, e.g. ImageNet, where $K$ is large and calculating a large matrix inverse is inherently time-consuming. Second, SimATTA employs a K-Means clustering algorithm for the selection of samples to annotate. Since K-Means is an iterative approach the computation cost of SimATTA is relatively high as well. In contrast, our proposed sample selection is built upon K-Center clustering which sequentially selects the sample with the highest distance to cluster centers (greedy algorithm), the computation cost is significantly lower than VeSSAL and SimATTA. An empirical study in Tab. 9 reveals that our method is at least 3 times faster than VeSSAL and SimATTA in Adaptation Time.
@@ B.5 Computation Complexity of Competing Methods
B.6 Studies for Regularization in HILTTA
We argue that proposed anchor regularization and EMA smoothing for model selection in HILTTA prevent the model from overly adapting to local data distribution on the testing data stream. Removing them may help adaptation to a single domain but harm the performance of future domains. As shown in Tab. 10, in the continual TTA setting, model selection without regularization is much worse than the one with regularization (17.25% v.s. 15.87%), suggesting regularization is necessary for continual TTA.
Regularization Gaussian Shot Impulse Defocus Jpeg AVG.
Without 25.04 20.23 28.57 12.22 19.37 17.25
With 25.93 20.13 27.49 11.59 18.26 15.87 (+1.38)
Table 10: Comparison between w/ and w/o regularization for HILTTA with TENT on Cifar10-C. error rates are reported.
B.7 Comparing with Knowledge Distillation and Ensemble Learning
We carried out additional experiments by comparing with ensemble learning and multi-teacher knowledge distillation methods. Specifically, for ensemble learning, we average the posterior predicted by each candidate model and make a prediction with the average posterior. For multi-teacher knowledge distillation, we use the pseudo label predicted by the ensemble model for self-training. As observed from Tab. 11, both ensemble learning and multi-teacher knowledge distillation yield inferior results than HILTTA. This is most likely caused by incorporating poor candidate models (ensemble learning) and using poor pseudo labels for self-training, i.e. learning from noisy labeled data (multi-teacher knowledge distillation).
Error 27.77 19.48 26.42 15.87
B.8 Random Batchsizes
We recognize that in certain scenarios, test data may arrive inconsistently and intermittently. Results from Tab. 12 demonstrate that our proposed HILTTA framework is adaptable to such varying batchsize. At each iteration, we vary the batchsize randomly among the choices [50, 100, 150, 200]. This versatility leads to even greater improvements compared to fixed batch size. This enhanced performance stems from the fact that the optimal hyperparameter values vary with different batch sizes, thus benefiting significantly from our adaptive model selection approach.
@@ B.8 Random Batchsizes
B.9 Ablation Study with RMT
We have conducted wider experiments to analyze the ablation study with RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] as the off-the-shelf TTA method in Tab. 13 exhibits a consistent trend with the findings presented in the manuscript, providing additional support for our analysis and demonstrating the effectiveness of our proposed approach.
@@ B.9 Ablation Study with RMT
Appendix C Experiment Details
C.1 Detailed Settings for Datasets
Following prior works [Döbler et al.(2023)Döbler, Marsden, and Yang, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan], we perform continual test time adaptation in the following datasets, detailed in follows.
CIFAR10-C [Hendrycks and Dietterich(2018)] is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 10 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
CIFAR100-C [Hendrycks and Dietterich(2018)] is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 100 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
ImageNet-C [Hendrycks and Dietterich(2018)] is a large-scale corruption dataset, with 15 different common corruptions, each containing 50,000 corrupt images of dimension (3, 224, 224) with 1000 categories. We evaluate the first 5,000 samples with severity level 5 following [Croce et al.(2021)Croce, Andriushchenko, Sehwag, Debenedetti, Flammarion, Chiang, Mittal, and Hein]. The domain sequence is "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg".
ImageNet-D [Rusak et al.(2022)Rusak, Schneider, Gehler, Bringmann, Brendel, and Bethge], as a large-scale style-transfer dataset, is built upon DomainNet [Peng et al.(2019)Peng, Bai, Xia, Huang, Saenko, and Wang], which encompasses 6 domain shifts (clipart, infograph, painting, quickdraw, real, and sketch). The dataset focuses on samples belonging to the 164 classes that overlap with ImageNet. Following [Marsden et al.(2024)Marsden, Döbler, and Yang], we specifically select all classes with a one-to-one mapping from DomainNet to ImageNet, resulting in a total of 109 classes and excluding the quickdraw domain due to the challenge of attributing many examples to a specific class. After processing, it contains 5 domain shifts with each 10,000 images selected in 109 classes. The domain sequence is "clipart $\rightarrow$ infograph $\rightarrow$ painting $\rightarrow$ real $\rightarrow$ sketch".
C.2 Detailed Settings for Comparing Methods
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell]. It focuses on updating batch normalization layers through entropy minimization. We follow the official implementation [^1] https://github.com/DequanWang/tent of TENT to update BN parameters.
PL [Lee et al.(2013)]. We implement self-supervised training on unlabeled data updating all parameters by cross-entropy with an entropy threshold $E=\theta\times log(\#class)$. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate for the rest.
SHOT [Liang et al.(2020)Liang, Hu, and Feng]. It freezes the linear classifier and trains the feature extractor by balancing prediction category distribution, coupled with pseudo-label-based self-training. We follow its official code [^2] https://github.com/tim-learn/SHOT to update the feature extractor but remove the threshold. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate.
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan]. It adapts batch normalization layers using entropy minimization but introduces an additional fisher regularization term to prevent drastic parameter changes. We follow the official implementation [^3] https://github.com/mr-eggplant/EATA of EATA to update BN parameters. We hold 2000 fisher training samples and keep the entropy threshold $E_{0}=0.4\times log(\#class)$.
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. It updates batch normalization layers with a sharpness-aware optimization approach. We mainly follow its official code [^4] https://github.com/mr-eggplant/SAR, but set $e_{0}=0$ for CIFAR10-C and CIFAR100-C, preventing failure cases of frequently resetting the model’s weight.
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang]. It adapts in a teacher-student mode, incorporating cross-entropy consistency loss and contrastive loss. We follow the official implementation [^5] https://github.com/mariodoebler/test-time-adaptation.
ASE [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth]. It uses surrogate estimators to sample in a uncertainty distribution to sample with low bias. We follow its official implementation [^6] https://github.com/jlko/active-surrogate-estimators.
SIMATTA [Gui et al.(2024)Gui, Li, and Ji]. As an active test-time adaptation method, it selects labeled samples through incremental clustering and updates the model by minimizing cross-entropy. To reduce the computational and memory cost of incremental clustering, we keep the maximum length of anchors to 50 in all experiments. For CIFAR10-C and CIFAR100-C, we set the lower entropy bound $e_{l}$ to 0.005 and 0.01 for the higher entropy bound $e_{u}$. While for ImageNet-C and ImageNet-D, we set the lower entropy bound $e_{l}$ to 0.2 and 0.4 for the higher entropy bound $e_{u}$.
VeSSAL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash]. As a stream-based active learning method, it uses volume sampling for streaming active learning. We also follow its official implementation [^7] https://github.com/asaran/VeSSAL, using feature embedding for stream-based data selection.
C.3 Hyper-parameter Choices for Model Selection
To achieve a fair and diverse validation process, we select three main kinds of hyper-parameters, i.e. threshold, loss coefficient, and learning rate. Detailed hyper-parameter settings along with the search space are available in Tab. 14.
TTA method Hyper-parameter Search Space
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
PL [Lee et al.(2013)] Entropy Threshold $\tau$0.05, 0.1, 0.2, 0.4, 0.6, 0.8, 1.0}
SHOT [Liang et al.(2020)Liang, Hu, and Feng] Loss Coefficient $\beta$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] Loss Coefficient $\lambda_{CL}$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}

### v2

@@ Appendix B Additional Experimental Analysis
Comparing Adaptation Time & Annotation Time: We carried out empirical studies by measuring the wall-clock time lapses of the inference step (Inference Time), adaptation step (Adaptation Time), and human annotation (Annotation Time) per sample. In specific, the "Inference Time" refers to the time-lapse of pure forward pass, the "Adaption Time" refers to the time-lapse of backpropagation and weights update, and "Annotation Time" refers to the estimated time lapse required for human to annotate images at a 3% labeling rate. We evaluate under three categories of annotation cost, i.e. 2 sec/sample, 5 sec/sample, and 10 sec/sample. As seen from Tab. 9, despite the adaptation time being higher than the inference time (8.7 ms v.s. 0.4 ms), the annotation time is the most expensive step. These results suggest model selection and adaptation do not slow down the throughput under the HILTTA protocol.
@@
Computation Complexity of Competing Active TTA Methods: We find that existing stream-based active learning (VeSSAL (Kossen et al., 2022)) and active TTA (SimATTA (Gui et al., 2024)) are inherently more computationally expensive than HILTTA due to the following reasons. First, VeSSAL adopts gradient embedding, i.e. $g(x_{i})=\partial l(f(x_{i};\theta),y_{i})/\partial\theta_{L}$, where $\theta_{L}\in\mathbb{R}^{K\times D}$ refers to the classifier weights. This results in a very high dimensional gradient embedding $g(x_{i})\in\mathbb{R}^{D\cdot K}$. The probability of labeling a sample $p_{i}$ involves calculating the inverse of covariance matrix, i.e. $p_{i}\propto g(x_{i})^{\top}\hat{\Sigma}^{-1}g(x_{i}),\;s.t.\hat{\Sigma}=\sum_{x_{i}\in\mathcal{B}}g(x_{i})g(x_{i})^{\top}\in\mathbb{R}^{DK\times DK}$. Hence, VeSSAL can hardly deal with classification tasks with a large number of classes, e.g. ImageNet, where $K$ is large and calculating a large matrix inverse is inherently time-consuming. Second, SimATTA employs a K-Means clustering algorithm for the selection of samples to annotate. Since K-Means is an iterative approach the computation cost of SimATTA is relatively high as well. In contrast, our proposed sample selection is built upon K-Center clustering which sequentially selects the sample with the highest distance to cluster centers (greedy algorithm), the computation cost is significantly lower. An empirical study in Tab. 10 reveals that our method is about 3 times faster than VeSSAL and SimATTA in adaptation time.
@@
Random Batchsizes: We recognize that in certain scenarios, test data may arrive inconsistently and intermittently. Results from Tab. 11 demonstrate that our proposed HILTTA framework is adaptable to such varying batchsize. At each iteration, we vary the batchsize randomly among the choices [50, 100, 150, 200]. This versatility leads to even greater improvements compared to fixed batch size. This enhanced performance stems from the fact that the optimal hyper-parameter values vary with different batch sizes, thus benefiting significantly from our adaptive model selection approach.
@@
Ablation Study with RMT: We have conducted wider experiments to analyze the ablation study with RMT (Döbler et al., 2023) as the off-the-shelf TTA method in Tab. 12 exhibits a consistent trend with the findings presented in the manuscript, providing additional support for our analysis and demonstrating the effectiveness of our proposed approach.